OpenStudy (anonymous):

A four digit number (with no repetitions) is to be formed from the set of numbers {0,1,2,3,4,5,6,7}. Find the probability that the number a) is even b) is odd c) is less than 4000 d) is less than 4000, given that it is greater than 3000

OpenStudy (dumbcow):

it should be somewhat obvious that it has a 50% chance of being either even or odd if it ends in even number, then its even if ends with odd, then its odd there are equal amount of odd/even numbers in set

hartnn (hartnn):

could u first find how many 4 digits number could be made ? no @dumbcow .

OpenStudy (dumbcow):

there are 8*7*6*5 possible numbers .... 1680 for each "thousand", there are 7*6*5 possible numbers ... 210 meaning 210 possible numbers between 0 and 999 210 possible numbers between 1000 and 1999 and so on ...

hartnn (hartnn):

part d is simplest: numbers greater than 3000 can start with 3, 4, 5, 6 or 7 numbers from the above sample space less than 4000 have to start with 3 reqd. probability = 1/5 = .2

hartnn (hartnn):

to form a 4 digit number,u cannot use those first 210 numbers in 0 to 999 so the total numbers that can be formed are 7*7*6*5=1470 or 1680-210=1470

hartnn (hartnn):

for part c.) the 1st digit can only be 1,2,3 for number to be <4000 the 2nd digit could be any of the remaining 7,3rd digit any of the remaining 6,4th digit any of the remaining 5 so the probability=3*7*6*5/7*7*6*5=3/7 got this?